# Algorithmic Primitives and Compositional Geometry of Reasoning in Language Models

### v1

@@ 5 Results
Primitive Tracing and Steering. We first traced primitives by clustering latent representations over entire reasoning traces, interpreting these clusters by identifying the corresponding tokens in the model output. Perhaps surprisingly, we found that despite the simplicity of our approach, many of these clusters were highly interpretable; for example, we discovered a cluster specifically associated with the implementation of a nearest-neighbor heuristic, and another cluster associated with validations, checks, and corrections.
We then determined the average transition probability between the different clusters. Spectral clustering of this transition matrix revealed that, on TSP, clusters are sequentially composed in a hierarchical and highly stereotyped manner, as most transitions between clusters occcurred within the same meta-cluster (Fig. 3a). Notably, these different meta-clusters were also responsible for highly interpretable parts of the reasoning, e.g. path generation or computing distances (Fig. 3c).
We then validated the role of these primitives by extracting the corresponding primitive vectors and injecting them into Phi-4 during a response generation to TSP. First, we injected two vectors (nearest_neighbor and generate_path) into the model throughout the response generation, exploring different layers and magnitudes. We found that nearest_neighbor selectively increased the proportion of nearest-neighbor paths generated by the model, whereas generate_path increased the number of total generated paths (Fig. 2a,b). This demonstrates that the extracted primitive vectors selectively induce a particular behavior and validates the candidate primitives generated by our clustering.
To expand this investigation, we injected a broader range of candidate primitive vectors into the model after providing example paths and distance computations. For example, injecting the compute_distance primitive causes the model to more quickly compute the distance of a candidate path. Interestingly, the reverse is also true: subtracting the compute_distance vector makes the model less likely to implement a distance computation (Fig. 2c). This suggests that subtracting primitive vectors can prevent associated algorithmic primitives. More broadly, we evaluate six behavioral hallmarks across six possible function vectors and find that all behavioral hallmarks are most strongly induced by a candidate primitive vector with a relevant computational role (Table 1).
@@
Table 1: Effects of primitive vector injection on behavioral hallmarks (% above baseline). For each cell, we identify the maximal effect across all positive magnitudes and all intervention layers (10, 13, 15, 17, 20, 30). Bold = strongest effect, underline = second strongest per column. (%NN paths: proportion of nearest neighbor paths generated. Dist. comp.: Distance computation. Final ans.: Final answer. #Verif.: number of Verifications in the output. #Comps: number of comparisons in the output. See Appendix D for detailed definitions.)
@@
Comparing Algorithmic Primitives between Phi-4 and Phi-4-Reasoning. To identify shared and distinct primitives between responses by Phi-4 and Phi-4-Reasoning, we computed the dissimilarity between the primitive frequencies involved in different responses. We fond that responses by Phi-4-Reasoning are highly stereotyped (Fig. 4a). In contrast, Phi-4 has two subgroups of distinct responses. Notably, these groups map onto two distinct approaches towards solving TSP: a brute-force search and a random guess without further refinement. This highlights that algorithmic fingerprinting reveals both differences between different models, but even differences within responses from the same model. Beyond TSP, we found that the responses from Phi-4 and Phi-4-Reasoning on 3-SAT and GraphNav are highly stereotyped (Fig. 4b,d). On AIME, algorithmic primitives involved in responses to the same question were very similar to each other, whereas primitives involved in responses to different questions were substantially more different, reflecting the higher diversity of questions on this benchmark (Fig. 4c).
@@
Next, we analyzed which primitives are more common in Phi-4 or Phi-4-Reasoning by computing the signed $\chi$-squared difference, and analyzing the clusters with the largest positive and negative values. This allows us to identify important differences in responses strategies across the two model. For example, on TSP and 3-SAT, distinct heuristic approaches arise more commonly in Phi-4-Reasoning than Phi-4 (nearest_neighbor and if_clause_true, Appendix A). Beyond task-specific strategies, our comparative analysis also highlights broader differences between the two models: in particular, clusters that are more common in Phi-4-Reasoning relate to general-purpose reasoning steps such as verification, recalling instructions, or planning the final answer.
Compositional Primitive Induction. So far, we have considered sequential compositions of algorithmic primitives. To investigate arithmetic compositions of primitives, we consider a set of algorithmic in-context learning tasks that require identifying 1) the last node of a path (“Terminal node recognition,” TNR), 2) the node with the higher reward (“Reward comparison,” RC); and 3) the most highly rewarded node between two paths, a composition of TNR and RC (Appendix C.1). Remarkably, we find that adding together the primitive vectors for TNR and RC induces this composite behavior in Llama-3-8B, causing it to match few-shot performance in a zero-shot setting (Fig. 5a). These effects were more attenuated in Phi-4 and Phi-4-Reasoning, highlighting that different models may require different modes of composition (Fig. 8).
Cross-task Transfer and Compositional Generalization. Generalized algorithmic primitives which can be applied across different tasks, can help us understand why reasoning-finetuned models like Phi-4-Reasoning show improvements on tasks they were not finetuned on. To better understand the degree to which algorithmic primitives in Phi-4-Reasoning are shared between tasks, we apply our algorithmic tracing framework to a set of responses from both TSP and AIME (Fig. 4e). This analysis highlights that while there are different algorithmic primitives at play in either task, the two tasks also have many algorithmic primitives in common.
Next, we tested whether cross-task injection of algorithmic primitives would induce behavioral changes. First, we extracted primitive vectors from an algorithmic in-context learning task that required extracting a path’s first or last node (Fig. 5b). We then injected these primitive vectors into complex reasoning traces solving TSP immediately after the model had mentioned a particular path. Despite the extremely different domain, injecting these vectors indeed had the expected effect: injecting the first_node (resp. last_node) primitive vector caused the model to subsequently mention the first node (resp. last node) of the path more often.
Finally, we considered relevant primitive vectors from AIME (e.g. spatial_reasoning, plan_final_answer) and injected them into Phi-4 generating a response to a TSP prompt. We found that these primitive vectors caused the model to generate more paths and a higher proportion of nearest-neighbor paths — a hallmark of Phi-4-Reasoning responses.
6 Discussion
Algorithmic Tracing and Steering. We introduce a framework for tracing and steering algorithmic primitives as building blocks of LLM reasoning with geometric compositionality. We adapted function vector techniques (Todd et al., 2024) to extract minimal, 1-step primitive vectors from carefully designed simple tasks as well as benchmarks with more complex reasoning contexts. When “injected” to LLM layers after the prompt or in the middle of reasoning traces, these primitive vectors increase the behavioral expression of corresponding reasoning steps (Figs. 2 and 5). This extends function vector research beyond simple relations between the input and output. A central contribution of our work is the compositional exploration of primitive vector arithmetic. Our framework bridges a principled path from laboratory-identified primitives to real-world reasoning behaviors.
@@ 6 Discussion
Toward a Geometry of Compositional Abstraction in LLMs. We find that algorithmic primitives exhibit geometric regularities through compositional operations like addition, subtraction, multiplication, and scalar modulation. The layer and magnitude of injection shape the output expression of the primitive (Fig. 2). We show cross-task transfer and generalizability of primitive vectors: spatial reasoning and verification primitives extracted from AIME successfully transfer to TSP. This transferability hints at potentially universal algorithmic building blocks underlying diverse reasoning capabilities. If this is the case, an algorithm may be expressed by an LLM in terms of an algebraic geometry of primitive vectors and their composition.
@@
Reasoning vs. Memorization. Our work contributes to recent discussions about the extent to which LLMs exhibit abstract reasoning rather than relying on memorized procedures, failing to generalize in counterfactual scenarios (Wu et al., 2024; Power et al., 2022; Zhang et al., 2024; Wang et al., 2025). While models were not directly trained on our tasks (Abdin et al., 2025), it is possible that they may be using amortized reasoning based on similar examples in the training. LLMs exhibit better problem-solving in high probability than low probability settings (e.g., McCoy et al. (2024a)), which rely on statistical regularities rather than abstract reasoning. Moreover, the absence of metacognitive abilities in LLMs results in brittle problem-solving (Johnson et al., 2024; Lewis & Mitchell, 2024). However, models finetuned for reasoning exhibit less sensitivity to task probability than base models (McCoy et al. (2024b)) and improved metacognitive-like uncertainty management strategies (e.g., verifying whether a candidate solution is correct) (Guo et al., 2025; Gandhi et al., 2025) that in turn causally improve reasoning (Bogdan et al., 2025). In line with these findings, our analysis identified algorithmic primitives for managing uncertainty (e.g., verification) and in-context recall. Such primitives and their associated behaviors may underlie a shift from memorization toward compositional problem-solving in reasoning models.
@@
Limitations. One limitation is that the primitives, corresponding clusters, and meta-clusters we identify do not always map to a known algorithm. Future work can extract primitive libraries and algorithmic logic of a given model, identifying commonalities and differences across models. Here, we focus on reasoning as multi-step problem-solving in limited tasks, which may not capture the complexity and multimodal nature of natural reasoning, studied in cognitive sciences for decades Tversky (2005); Shepard & Metzler (1971); MacGregor & Ormerod (1996). Finally, we have focused on linear compositions of primitive vectors, leaving more complex interactions and potentially non-linear combinations of several vectors on manifolds for future work.
@@
Future Directions. The algorithmic tracing and steering framework can be applied to any architecture (e.g., vision, diffusion, and multi-modal models) and domain beyond reasoning. An immediate extension is more detailed manifold analysis in order to capture the compositional geometry of primitives beyond linear composition, and establish universal and task-specific primitive libraries. Another key direction is the algorithmic training and finetuning of LLMs, with algorithmic objectives. Moreover, future self-improving models can be designed for algorithmic self-play: generating and evaluating compositional algorithmic solutions. Finally, collecting human reasoning data on the same tasks enables us to compare and finetune model-human algorithmic alignment.
@@
LLM Usage Disclosure
We used LLMs for generating code and for finding related work.
@@ A.1 Detailed methods
In all cases, we used extracted the token-by-token representation from layer 17 of Phi-4-base. We used k-means clustering with $k=50$ clusters and designed a html interface to visually inspect the different clusters.
@@ A.2 Phi-4-Reasoning responses on TSP
Cluster 37: nearest_neighbor. This cluster is active specifically during the tokens where the model identifies the nearest-neighbor path or the closest city within a particular candidate path. As Phi-4-Reasoning’s responses, in contrast to those by Phi-4, often started out by an initial guess using the nearest-neighbor heuristic (see Appendix ??), we hypothesized that this would be a particularly relevant primitive.
@@ A.3.1 Clusters more common in Phi-4-Reasoning
A.4 Responses by Phi-4 and Phi-4-Reasoning on 3SAT
Clusters more common in Phi-4-Reasoning.
@@ A.4 Responses by Phi-4 and Phi-4-Reasoning on 3SAT
A.4.1 Clusters more common in Phi-4.
@@
A.4.2 Clusters common to both models.
@@ B.2 American Invitational Mathematics Examination (AIME)
B.3 3-literal Satisfiability Problem (3SAT)
Models were tasked with determining the satisfiability of various 3SAT problems, a class of NP-hard problems requiring combinatorial reasoning. If a model responded that a problem was satisfiable, they were required to provide the speoicific literals that would achieve satisfiability.

### v2

@@ 5 Results
Table 2: Effects of primitive vector injection on behavioral hallmarks (% above baseline). For each cell, we identify the maximal effect across all positive magnitudes and intervention layers (10, 13, 15, 17, 20, 30). Bold = strongest effect, underline = second strongest per column. (%NN paths: Proportion of nearest neighbor paths generated. Dist. comp.: Distance computation. Final ans.: Final answer. #Verif.: number of Verifications in the output. #Comps: number of comparisons in the output. See App. F for detailed definitions.)
@@
Primitive Tracing and Steering. We first traced primitives by clustering latent representations over entire reasoning traces, interpreting these clusters by identifying the corresponding tokens in the model output. Perhaps surprisingly, we found that despite the simplicity of our approach, many of these clusters were highly interpretable; for example, we discovered a cluster specifically associated with the implementation of a nearest-neighbor heuristic, and another cluster associated with validations, checks, and corrections. Notably, we found that the labels provided by the LLM-driven annotation pipeline were highly consistent with our manual annotations. This validates the accuracy of the LLM-driven pipeline (Table LABEL:tab:tsp).
We then determined the average transition probability between the different clusters. Spectral clustering of this transition matrix revealed that, on TSP, clusters are sequentially composed in a hierarchical and highly stereotyped manner, as most transitions between clusters occurred within the same meta-cluster (Fig. 2a). Notably, different meta-clusters were also responsible for highly interpretable parts of the reasoning, e.g. path generation or computing distances (Fig. 2c).
Next, we validated the role of these primitives by extracting the corresponding primitive vectors and injecting them into Phi-4 during a response generation to TSP. First, we injected two vectors (nearest_neighbor and generate_path) into the model throughout the response generation, exploring different layers and magnitudes. We found that nearest_neighbor selectively increased the proportion of nearest-neighbor paths generated by the model, whereas generate_path increased the number of total generated paths (Fig. 3a,b). This demonstrates that the extracted primitive vectors selectively induce a particular behavior and validates the candidate primitives generated by our clustering.
To expand this investigation, we injected a broader range of candidate primitive vectors into the model after providing example paths and distance computations. For example, injecting the compute_distance primitive causes the model to more quickly compute the distance of a candidate path. Interestingly, the reverse is also true: subtracting the primitive makes the model less likely to implement a distance computation (Fig. 3c). This suggests that subtracting primitive vectors can prevent associated algorithmic primitives. More broadly, we evaluate six behavioral hallmarks across six possible function vectors and find that all behavioral hallmarks are most strongly induced by a candidate primitive vector with a relevant computational role (Table 2; see further control experiments in App. F.1, F.3).
@@
Comparing Algorithmic Primitives between Models. To compare primitives between Phi-4 and Phi-4-Reasoning, we computed the dissimilarity between the primitive frequencies involved in their responses. We found that responses by Phi-4-Reasoning are highly stereotyped (Fig. 4a). In contrast, Phi-4 has two subgroups of distinct responses. Notably, these groups map onto two distinct approaches towards solving TSP: a brute-force search and a random guess without further refinement. This highlights that algorithmic fingerprinting reveals both differences between models, and differences between responses from the same model. Beyond TSP, the responses from Phi-4 and Phi-4-Reasoning on 3-SAT and GraphNav are highly stereotyped (Fig. 4b,d). On AIME, algorithmic primitives involved in responses to the same question were very similar to each other, whereas primitives involved in responses to different questions were substantially more different, reflecting the higher diversity of questions in AIME (Fig. 4c).
Next, we analyzed which primitives are more common in Phi-4 or Phi-4-Reasoning by computing the signed $\chi$-squared difference, and analyzing the clusters with the largest positive and negative values. This allows us to identify important differences in response strategies across the two models. For example, on TSP and 3-SAT, distinct heuristic approaches arise more commonly in Phi-4-Reasoning than Phi-4 (nearest_neighbor and if_clause_true, App. A). Beyond task-specific strategies, our comparative analysis also highlights broader differences between the two models: in particular, clusters that are more common in Phi-4-Reasoning relate to general-purpose reasoning steps such as verification, recalling instructions, or planning the final answer.
Compositional Primitive Induction. So far, we have considered sequential compositions of algorithmic primitives. To investigate arithmetic compositions of primitives, we consider a set of algorithmic in-context learning tasks that require identifying 1) the last node of a path (“Terminal node recognition,” TNR), 2) the node with the higher reward (“Reward comparison,” RC; reward was defined by a number between 1 and 100); and 3) the most highly rewarded node between two paths, a composition of TNR and RC (see example prompt in App. E.1). Remarkably, we find that adding together the primitive vectors for TNR and RC induces this composite behavior in Llama-3-8B, causing it to match few-shot performance in a zero-shot setting (Fig. 5a). These effects were more attenuated in Phi-4 and Phi-4-Reasoning, highlighting that different models may require different modes of composition (Fig. 10).
Cross-task and Compositional Generalization. Generalized algorithmic primitives which can be applied across different tasks, can help us understand why reasoning-finetuned models like Phi-4-Reasoning show improvements on tasks they were not finetuned on. To better understand the degree to which algorithmic primitives in Phi-4-Reasoning are shared between tasks, we apply our algorithmic tracing framework to a set of responses from both TSP and AIME (Fig. 4e). This analysis highlights that while there are different algorithmic primitives at play in either task, the tasks also have many primitives in common.
Next, we tested whether cross-task injection of algorithmic primitives would induce behavioral changes. We extracted primitive vectors from an algorithmic in-context learning task that required extracting a path’s first or last node (Fig. 5b) and injected them into complex reasoning traces solving TSP immediately after the model had mentioned a particular path. Despite the extremely different domain, injecting these vectors indeed had the expected effect: injecting the get_first_node (resp. get_last_node) primitive vector caused the model to mention the first node (resp. last node) of the path more often (Fig. 5c).
Finally, we considered relevant primitive vectors from AIME (e.g. spatial_reasoning, plan_final_answer) and injected them into Phi-4 generating a response to a TSP prompt. We found that these primitive vectors caused the model to generate more paths and a higher proportion of nearest-neighbor paths (Fig. 3d) — a hallmark of Phi-4-Reasoning responses.
6 Discussion
Algorithmic Tracing and Steering. We introduce a framework for tracing and steering algorithmic primitives as building blocks of LLM reasoning with geometric compositionality. We adapted function vector techniques (Todd et al., 2024) to extract minimal, 1-step primitive vectors from carefully designed simple tasks as well as benchmarks with more complex reasoning contexts. When “injected” to LLM layers after the prompt or in the middle of reasoning traces, these primitive vectors increase the behavioral expression of corresponding reasoning steps (Figs. 3 and 5). This extends function vector research beyond simple input-output relations. A central contribution of our work is the compositional exploration of primitive vector arithmetic. Our framework bridges a principled path from laboratory-identified primitives to real-world reasoning behaviors.
@@ 6 Discussion
Toward a Geometry of Compositional Abstraction in LLMs. We find that algorithmic primitives exhibit geometric regularities through compositional operations like addition, subtraction, multiplication, and scalar modulation. The layer and magnitude of injection shape the expression of the primitive (Fig. 3). We show cross-task transfer and generalizability of primitive vectors: spatial reasoning and verification primitives extracted from AIME successfully transfer to TSP. This transferability hints at potentially universal algorithmic building blocks underlying diverse reasoning capabilities. If this is the case, an algorithm may be expressed by an LLM in terms of an algebraic geometry of primitive vectors and their composition.
@@
Reasoning vs. Memorization. Our work contributes to recent discussions about the extent to which LLMs exhibit abstract reasoning rather than relying on memorized procedures, failing to generalize in counterfactual scenarios (Wu et al., 2024; Power et al., 2022; Zhang et al., 2024; Wang et al., 2025). While models were not directly trained on our tasks (Abdin et al., 2025), it is possible that they may be using amortized reasoning based on similar examples in the training. LLMs exhibit better problem-solving in high probability than low probability settings (e.g. McCoy et al., 2024a), which rely on statistical regularities rather than abstract reasoning. Moreover, the absence of metacognitive abilities in LLMs results in brittle problem-solving (Johnson et al., 2024; Lewis and Mitchell, 2024). However, models finetuned for reasoning exhibit less sensitivity to task probability than base models McCoy et al. (2024b) and improved metacognitive-like uncertainty management strategies (e.g., verifying whether a candidate solution is correct) (Guo et al., 2025; Gandhi et al., 2025) that in turn causally improve reasoning (Bogdan et al., 2025). In line with these findings, our analysis identified algorithmic primitives for managing uncertainty (e.g., verification) and in-context recall. Such primitives and their associated behaviors may underlie a shift from memorization toward compositional problem-solving in reasoning models.
@@
Limitations. One limitation is that the primitives, corresponding clusters, and meta-clusters we identify do not always map to a known algorithm. Future work can identify primitive ontology and algorithmic logic of different models, identifying commonalities and differences across them. Here, we focus on reasoning as multi-step problem-solving in limited tasks, which may not capture the complexity and multi-modal nature of natural reasoning, studied in cognitive sciences for decades Tversky (2005); Shepard and Metzler (1971); MacGregor and Ormerod (1996). Finally, we have focused on linear compositions of primitive vectors, leaving more complex interactions, e.g. non-linear combinations of several vectors on manifolds, for future work.
@@
Future Directions. The algorithmic tracing and steering framework can be applied to any architecture (e.g., vision, diffusion, and multi-modal models) and domain beyond reasoning. An immediate extension is a more detailed manifold analysis in order to capture the compositional geometry of primitives beyond linear composition, and establish a task-specific and universal primitive ontology. Another key direction is the algorithmic training and finetuning of LLMs, with algorithmic objectives. Future self-improving models can be designed for algorithmic self-play: generating and evaluating compositional algorithmic solutions. Finally, collecting human reasoning data on the same tasks enables us to compare and finetune model-human algorithmic alignment.
@@ Acknowledgments
Impact Statement
Our work offers a rigorous and multi-level framework for the evaluation and understanding of reasoning behavior in large language models (LLMs). Current LLMs are increasingly capable, but can fail in unpredictable ways at reasoning and can display harmful biases (Gallegos et al., 2024). Our approach offers an algorithmic understanding of LLM reasoning at both the behavioral level of reasoning traces and the internal level of latent processes and layer by layer representations. Such an algorithmic understanding can help us understand the contexts in which LLMs’ algorithmic primitives can implement correct reasoning behavior and the contexts in which we would expect them to make mistakes. This improved understanding, in turn, would allow us to avoid deploying these models in contexts where we would expect them to make mistakes. In addition, our approach could flag the specific algorithmic primitives that give rise to particular biases, and offer a path to target finetuning to those specific algorithmic failures. Finally, our approach highlights the fundamental steps required to solve a reasoning problem, offering empirical ways to measure and compare the sequence of primitives generated by the LLM. In the future, this research can inspire improved and more sample-efficient training procedures as well as novel architectures. The sample efficiency of architectures trained with an algorithmic objective of compositional generalization can advance artificial intelligence in two ways, both improving the quality of performance and reducing the ecological impact of training. Such advances have many potential societal consequences, consistent with foundational research on machine learning models more generally.
@@ A.1 Detailed methods
In all cases, we extracted the token-by-token representation from layer 17 of Phi-4-base. We used k-means clustering with $k=50$ clusters and designed an HTML interface to visually inspect the different clusters (we will make this tool available on the project website).
@@ A.2 Phi-4-Reasoning responses on TSP
Cluster 37: nearest_neighbor. This cluster is active specifically during the tokens where the model identifies the nearest-neighbor path or the closest city within a particular candidate path. As Phi-4-Reasoning’s responses, in contrast to those by Phi-4, often started out by an initial guess using the nearest-neighbor heuristic (see Appendix D), we hypothesized that this would be a particularly relevant primitive.
@@ A.3.1 Clusters more common in Phi-4-Reasoning
A.4 Responses by Phi-4 and Phi-4-Reasoning on 3-SAT
A.4.1 Clusters more common in Phi-4-Reasoning.
@@ A.4 Responses by Phi-4 and Phi-4-Reasoning on 3-SAT
A.4.2 Clusters more common in Phi-4.
@@
A.4.3 Clusters common to both models.
@@ A.6 Responses by Phi-4 and Phi-4-Reasoning on GraphNav
A.7 Cluster expressivity
To ease the automated evaluation of identified clusters, we introduce a novel metric to quantify the expressivity of each cluster. “Cluster expressivity” measures the number of unique tokens expressed by a cluster, divided by the radius of that cluster in latent space. More formally, consider a response with tokens $T\in\mathcal{T}^{n}$ where $\mathcal{T}$ is the set of tokens. Let $x\in\mathbb{R}^{n\times d}$ be its latent representation and $C\in\mathcal{C}^{n}$ be its associated cluster identities. For each $c\in\mathcal{C}$, let $\text{cent}[c]\in\mathbb{R}^{d}$ be the associated centroid. We then define the radius
$\text{radius}[c]:=\max_{n:C_{n}=c}\|x_{n}-\text{cent}[c]\|_{2}$ (3)
as the maximal $\ell_{2}$-distance of a data point within this cluster. We define the number of unique tokens expressed by the cluster as
$m[c]:=|\{t\in\mathcal{T}:\exists_{n}C_{n}=c,T_{n}=t\}|.$ (4)
We then compute “cluster expressivity” as
$\frac{\text{radius}[c]}{m[c]}.$ (5)
To summarize this metric, we compute its median value across all responses. We reason that high expressivity indicates better-defined operations with rich vocabulary, i.e. multiple valid ways to express the same operation. We therefore take them as more likely to be accurate primitives with clear algorithmic meaning. Table LABEL:tab:tsp shows metrics of expressivity for all clusters extracted from the responses by Phi-4-Reasoning on TSP. Notably, our selected primitives consistantly have high cluster expressivity.
@@ B.2 American Invitational Mathematics Examination (AIME)
B.3 3-literal Satisfiability Problem (3-SAT)
Models were tasked with determining the satisfiability of various 3-SAT problems, a class of NP-hard problems requiring combinatorial reasoning. If a model responded that a problem was satisfiable, they were required to provide the specific literals that would achieve satisfiability.